Two anglesare complementary, and one is twice as large as the other one.cWhat is the measure of each angle?
let x= one angle let y= one angle
Would my equations be:
x+y=90
2x=y
I am not getting a correct answer, and my I believe the problem is my equations.

Answer by ankor@dixie-net.com(22740)

Two angles are complementary, and one is twice as large as the other one.c What is the measure of each angle?
let x= one angle let y= one angle
Would my equations be:
x+y=90
2x=y
This is right, now replace y with 2x in the 1st equation
x + 2x = 90
3x = 90
x = 90/3
x = 30 degrees
then obviously
y = 2x
y = 2(30)
y = 60 degrees
:
Are they complementary? 30 + 60 = 90